# Production of Dark Photons through Higher Electromagnetic Moments at LDMX: Simulations and Model Discrimination

### v1

@@ 4 Signatures at LDMX
LDMX [14], a missing momentum experiment which uses an electron beam incident on a tungsten target, will have the ability to reconstruct both the energy, $E_{e^{-}_{f}}$, and the transverse momentum (momentum component perpendicular to the beam), $|p_{T}|_{e^{-}_{f}}$, of the recoil electrons. The events are expected to have different kinematic distributions whether or not there has been DM or dark mediator production, as opposed to the production of ordinary photons, since the production of an invisible but massive particle will carry away most of the energy and kick the recoil electron in the transverse direction. As we will show in this section, the nature of the DM signal also can vary the shape of the kinematic distributions and total expected number of events. Thus, the kinematic distributions will vary depending on the dark photon model, namely the type of interaction between the dark photon and SM fermions. The expected kinematic distributions are calculated analytically and numerically, with disagreement due to assumptions made in the analytic approach. We then proceed with the numerical approach, and compute the projected sensitivity in reference to our calculated thermal targets.
@@ 4.1 Analytic vs Numerical Methods
Using the formalism presented in [20], which relies on the Weizsacker-Williams Approximation [53, 54], and comparing with the results of [23], we calculate the $E_{e^{-}_{f}}$ and $|p_{T}|_{e^{-}_{f}}$ distributions, namely $\frac{d\sigma}{dx_{A^{\prime}}d\cos\theta}$, in the kinetic mixing and magnetic dipole models. The differential cross section for dark photon production in the kinetic mixing model is given by [20],
$\frac{d\sigma_{KM}}{dx_{A^{\prime}}d\cos\theta}=Ax_{A^{\prime}}\left[\frac{1-x_{A^{\prime}}+x_{A^{\prime}}^{2}/2}{U^{2}}+\frac{(1-x_{A^{\prime}})^{2}m_{A^{\prime}}^{2}}{U^{4}}\left(m_{A^{\prime}}^{2}-\frac{Ux_{A^{\prime}}}{1-x_{A^{\prime}}}\right)\right],$ (4.1)
where $A$ is a constant which depends on $\epsilon$, $E_{\text{beam}}$, an effective flux of photons depending on the nuclear structure, and $m_{A^{\prime}}$, $x_{A^{\prime}}=E_{A^{\prime}}/E_{\text{beam}}$, $\theta$ is the angle of the dark photon with the beam axis, and $U=E_{\text{beam}}^{2}\theta^{2}x+m_{A^{\prime}}^{2}\frac{(1-x)}{x}+m_{e}^{2}x$. Similarly, from [23] we have the differential cross section for dark photon production in the magnetic dipole model,
$\frac{d\sigma_{M}}{dx_{A^{\prime}}d\cos\theta}=\frac{Bx_{A^{\prime}}^{2}}{U},$ (4.2)
where $B$ is a constant. The distributions of the recoil electron are computed by integrating $\frac{d\sigma}{dx_{A^{\prime}}d\cos\theta}$ over $\theta$ from 0 to $\pi$, and normalizing to properly compare with the numerical method described below. These analytic expressions rely on a number of assumptions. The dark photon is assumed to be roughly collinear with the beam, hence $\theta$ is small. In addition, the nucleus is assumed to remain at rest i.e. not recoil at all during this process.
Keeping these assumptions of the analytic method in mind, we simulate this dark bremsstrahlung process by running MadGraph [30] with our FeynRules [55] generated UFO files for each model, which include the nucleus-photon vertex with a form factor [20].
@@
The $p_{T}$ distribution of the final state tungsten nucleus, directly after the process, is plotted in Fig. 3 using MadGraph. There are a significant number of events where the nucleus receives $>100\,\mathrm{MeV}$ transverse momenta, a non-negligible quantity not accounted for in the analytic method. The momentum transfer to the nucleus is also neglected in the software packages Geant4 [56] and EGS5 [57] for modeling bremsstrahlung in the SM. In [58], the authors similarly point out problems with this assumption. For this reason, we continue this section using MadGraph to simulate dark bremsstrahlung events for each model, computing the projected sensitivity and expected kinematic distributions potentially used for model discrimination in case of a future signal at LDMX.
@@ 4.3 Kinematic Distributions
As mentioned above, and shown in [23], different models lead to different distributions of the recoil electron at missing momentum experiments, such as LDMX.
@@
For three representative dark photon masses, the transverse momentum and the energy of the final state electron, $e^{-}_{f}$, are plotted in Fig. 4 for each model. Similarly to the projected sensitivity curves, Fig. 1, the magnetic and electric dipole models are qualitatively similar to eachother, as well as the charge radius and anapole moment models. However, the kinetic mixing model distributions have a shape similar to the distributions of charge radius and anapole moment: from Eq. 2.1, since $q^{2}=m_{A^{\prime}}^{2}$ in the on-shell dark photon bremsstrahlung, there is no longer momentum dependence in the Feynman rule for the charge radius and anapole models, just as for kinetic mixing. The overall scaling of the kinetic mixing model events would be much larger, however the distributions have been normalized - bringing the number of groups from three to two. The difference in shapes between the two groups arises from the momentum dependence of the Feynman rules. The charge and anapole models have momentum dependent vertices, where the Feynman rule is $\propto q$, the momentum of the dark photon.
@@ 4.4 Degeneracy of Mass and Model
With enough signal events, the dark photon mass can be estimated based on the shape of the kinematic distributions [14, 15]. Interestingly, considering higher order electromagnetic moment interactions in addition to the standard kinetic mixing, complicates the mass estimation process, since varying the model also alters the kinematic distributions. Increasing $m_{A^{\prime}}$ pushes the $E_{e^{-}_{f}}$ distribution closer to zero and the $|p_{T}|_{e^{-}_{f}}$ to larger momenta, and on the other hand the magnetic and electric dipole moment models give rise to more higher energy and $p_{T}$ recoil electrons. Therefore, varying the model can have indistinguishable effects from varying the mass.
@@ 5 Conclusion
We consider an extended theoretical framework for the dark photon, where there are effective interactions involving loops of portal matter fields, in addition to the well studied electromagnetic-like kinetic mixing interaction, as presented in [23]. Due to the different Lorentz structures of the $A^{\prime}-f$ interaction vertices leading to different energy dependencies in the dark bremsstrahlung and DM annihilation cross sections, the expected signal distributions of LDMX and relic targets vary between models. By studying each new type of interaction separately, we compute the kinematic distributions expected at LDMX, the projected sensitivities of LDMX, and the relic targets – extending the theory reach of LDMX to include these additional models. Implications for other future and current experiments is reserved for future studies. We find that the models can be classified into three groups, each characterized by similar vertices thus quantitatively similar expected LDMX signals and relic targets: the first with the kinetic mixing model, the second with the magnetic and electric dipole models, and the third with the charge radius and anapole moment models.
@@ Appendix A Matrix Elements
[…]
@@
[…]
@@
$\overline{|\mathcal{M}|^{2}}_{\text{KM}}=\frac{2g_{D}^{2}\epsilon^{2}e^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}^{2}})\left[s-(s-4m_{f}^{2})\cos^{2}\theta\right]$ (A.3)
$\overline{|\mathcal{M}|^{2}}_{\text{M}}=\frac{4sg_{D}^{2}\mu_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})\left[(s-4m_{f}^{2})\cos^{2}\theta+4m_{f}^{2}\right]$ (A.4)
$\overline{|\mathcal{M}|^{2}}_{\text{E}}=\frac{4sg_{D}^{2}d_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})(s-4m_{f}^{2})\cos^{2}\theta$ (A.5)
$\overline{|\mathcal{M}|^{2}}_{\text{C}}=\frac{4s^{2}g_{D}^{2}b_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})\left[s-(s-4m_{f}^{2})\cos^{2}\theta\right]$ (A.6)
$\overline{|\mathcal{M}|^{2}}_{\text{A}}=\frac{4s^{2}g_{D}^{2}a_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})(s-4m_{f}^{2})\sin^{2}\theta$ (A.7)
@@
$\overline{|\mathcal{M}_{A^{\prime}\to\chi\overline{\chi}}|^{2}}=\frac{g_{D}^{2}}{3}(m_{A^{\prime}}^{2}-4m_{\text{DM}}^{2})$ (A.8)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{KM}}=\frac{e^{2}\epsilon^{2}}{3}(4m_{A^{\prime}}^{2}+8m_{f}^{2})$ (A.9)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{M}}=\frac{\mu_{f}^{2}}{3}(2m_{A^{\prime}}^{4}+16m_{f}^{2}m_{A^{\prime}}^{2})$ (A.10)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{E}}=\frac{d_{f}^{2}}{3}(m_{A^{\prime}}^{4}-4m_{f}^{2}m_{A^{\prime}}^{2})$ (A.11)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{C}}=\frac{b_{f}^{2}}{3}(4m_{A^{\prime}}^{6}+8m_{f}^{2}m_{A^{\prime}}^{4})$ (A.12)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{A}}=\frac{a_{f}^{2}}{3}(4m_{A^{\prime}}^{6}-16m_{f}^{2}m_{A^{\prime}}^{4})$ (A.13)

### v2

@@ 4 Signatures at LDMX
LDMX LDMX:2018cma , a missing momentum experiment which uses an electron beam incident on a tungsten target, will have the ability to reconstruct both the energy, $E_{e^{-}_{f}}$, and the transverse momentum (momentum component perpendicular to the beam), $|p_{T}|_{e^{-}_{f}}$, of the recoil electrons. The events are expected to have different kinematic distributions whether or not there has been DM or dark mediator production, as opposed to the production of ordinary photons, since the production of an invisible but massive particle will carry away most of the energy and kick the recoil electron in the transverse direction. As we will show in this section, the nature of the DM signal also can vary the shape of the kinematic distributions and total expected number of events. Thus, the kinematic distributions will vary depending on the dark photon model, namely the type of interaction between the dark photon and SM fermions. The expected kinematic distributions are calculated analytically and numerically, with disagreement due to assumptions made in the analytic approach. We then proceed with the numerical approach, and compute the projected sensitivity in reference to our calculated thermal targets.
@@ 4.1 Analytic vs Numerical Methods
Using the formalism presented in Bjorken:2009mm , which relies on the Weizsacker-Williams Approximation vonWeizsacker:1934nji ; Williams:1935dka , and comparing with the results of Rizzo:2021lob , we calculate the $E_{e^{-}_{f}}$ and $|p_{T}|_{e^{-}_{f}}$ distributions, namely $\frac{d\sigma}{dx_{A^{\prime}}d\cos\theta}$, in the kinetic mixing and magnetic dipole models. The differential cross section for dark photon production in the kinetic mixing model is given by Bjorken:2009mm ,
$\frac{d\sigma_{KM}}{dx_{A^{\prime}}d\cos\theta}=Ax_{A^{\prime}}\left[\frac{1-x_{A^{\prime}}+x_{A^{\prime}}^{2}/2}{U^{2}}+\frac{(1-x_{A^{\prime}})^{2}m_{A^{\prime}}^{2}}{U^{4}}\left(m_{A^{\prime}}^{2}-\frac{Ux_{A^{\prime}}}{1-x_{A^{\prime}}}\right)\right],$ (7)
where $A$ is a constant which depends on $\epsilon$, $E_{\text{beam}}$, an effective flux of photons depending on the nuclear structure, and $m_{A^{\prime}}$, $x_{A^{\prime}}=E_{A^{\prime}}/E_{\text{beam}}$, $\theta$ is the angle of the dark photon with the beam axis, and $U=E_{\text{beam}}^{2}\theta^{2}x+m_{A^{\prime}}^{2}\frac{(1-x)}{x}+m_{e}^{2}x$. Similarly, from Rizzo:2021lob we have the differential cross section for dark photon production in the magnetic dipole model,
$\frac{d\sigma_{M}}{dx_{A^{\prime}}d\cos\theta}=\frac{Bx_{A^{\prime}}^{2}}{U},$ (8)
where $B$ is a constant. Notice that the differential cross sections do not depend on parameters associated with the DM particle – only the dark photon mass and coupling to electrons. The distributions of the recoil electron are computed by integrating $\frac{d\sigma}{dx_{A^{\prime}}d\cos\theta}$ over $\theta$ from 0 to $\pi$, and normalizing to properly compare with the numerical method described below. These analytic expressions rely on a number of assumptions. The dark photon is assumed to be roughly collinear with the beam, hence $\theta$ is small. In addition, the nucleus is assumed to remain at rest i.e. not recoil at all during this process.
Keeping these assumptions of the analytic method in mind, we simulate this dark bremsstrahlung process by running MadGraph Alwall:2014hca with our FeynRules Alloul:2013bka generated UFO files for each model. We include the form factors for the nucleus-photon vertex given in A18 and A19 of Bjorken:2009mm  [^5] As pointed out by , there is a typo in eq. A19, where the second term should not be squared., which include both elastic and inelastic contributions. The signal contribution from a dark photon emitted off of the nucleus rather than the electron is safely neglected, since we expect the overall number of events and generated $p_{T}$ to both be small, however we leave a detailed study of the contribution to future work.
@@
The $p_{T}$ distribution of the final state tungsten nucleus, directly after the process, is plotted in Fig. 3 using MadGraph. There are a significant number of events where the nucleus receives $100\,\mathrm{MeV}$ transverse momenta, a non-negligible quantity not accounted for in the analytic method. The momentum transfer to the nucleus is also neglected in the software packages Geant4 GEANT4:2002zbu and EGS5 Hirayama:2005zm for modeling bremsstrahlung in the SM. In Blinov:2024pza , the authors similarly point out problems with this assumption. For this reason, we continue this section using MadGraph to simulate dark bremsstrahlung events for each model, computing the projected sensitivity and expected kinematic distributions potentially used for model discrimination in case of a future signal at LDMX.
@@ 4.3 Kinematic Distributions
As mentioned above, and shown in Rizzo:2021lob , different models lead to different distributions of the recoil electron at missing momentum experiments, such as LDMX.
@@
For three representative dark photon masses, the transverse momentum and the energy of the final state electron, $e^{-}_{f}$, are plotted in Fig. 4 for each model. Similarly to the projected sensitivity curves, Fig. 1, the magnetic and electric dipole models are qualitatively similar to eachother, as well as the charge radius and anapole moment models. However, the kinetic mixing model distributions have a shape similar to the distributions of charge radius and anapole moment: from Eq. 1, since $q^{2}=m_{A^{\prime}}^{2}$ in the on-shell dark photon bremsstrahlung, there is no longer momentum dependence in the Feynman rule for the charge radius and anapole models, just as for kinetic mixing. The overall scaling of the kinetic mixing model events would be much larger, however the distributions have been normalized - bringing the number of groups from three to two. The difference in shapes between the two groups arises from the momentum dependence of the Feynman rules. The charge and anapole models have momentum dependent vertices, where the Feynman rule is dependent on $q$, the momentum of the dark photon.
Figure 5: The normalized kinematic distributions of Fig. 4, plotted as 2-dimensional histograms in the $|p_{T}|_{e^{-}_{f}}$ vs $E_{e^{-}_{f}}$ plane. The kinetic mixing, charge radius, and anapole moment models are plotted on the left, while the magnetic and electric dipole models are on the right; where the two groupings are chosen since the distributions are the same within them apart from statistical fluctuations. The dark photon mass is set to $m_{A^{\prime}}=1\,\mathrm{GeV}$.
Fig. 5 plots the distributions of Fig. 4, but as a 2-dimensional histogram, highlighting the correlations between the two kinematic parameters measurable at LDMX. As evident by Fig. 4 and already discussed above, the models can be grouped by their signal behaviour in which their distributions are the same apart from statistical fluctuations. Therefore, since the distributions for the models belonging to each group are the same, neglecting statistical fluctuations, the distributions are presented as one plot per group. The left and right 2-dimensional distributions indeed do differ; there are more events towards higher energy and $p_{T}$ for the magnetic and electric dipole models. All events obey $E_{e_{f}^{-}}>|p_{T}|_{e_{f}^{-}}$, as expected, and most have energy and $p_{T}$ $\sim 1\,\mathrm{GeV}$.
@@ 4.4 Degeneracy of Mass and Model
Considering the ordinary kinetic mixing dark photon model with enough signal events, the dark photon mass can be estimated based on the shape of the kinematic distributions LDMX:2018cma ; Akesson:2022vza . This is possible since the only model parameter altering the shape of the distributions is $m_{A^{\prime}}$. Interestingly, considering higher order electromagnetic moment interactions in addition to the standard kinetic mixing, complicates the mass estimation process, since varying the model also alters the kinematic distributions. Increasing $m_{A^{\prime}}$ pushes the $E_{e^{-}_{f}}$ distribution closer to zero and the $|p_{T}|_{e^{-}_{f}}$ to larger momenta, and on the other hand the magnetic and electric dipole moment models give rise to more higher energy and $p_{T}$ recoil electrons. Therefore, varying the model can have indistinguishable effects from varying the mass.
@@ 5 Conclusion
We consider an extended theoretical framework for the dark photon, where there are effective interactions involving loops of portal matter fields, in addition to the well studied electromagnetic-like kinetic mixing interaction, as presented in Rizzo:2021lob . Due to the different Lorentz structures of the $A^{\prime}-f$ interaction vertices leading to different energy dependencies in the dark bremsstrahlung and DM annihilation cross sections, the expected signal distributions of LDMX and relic targets vary between models. By studying each new type of interaction separately, we compute the kinematic distributions expected at LDMX, the projected sensitivities of LDMX, and the relic targets – extending the theory reach of LDMX to include these additional models. Implications for other future and current experiments is reserved for future studies. We find that the models can be classified into three groups, each characterized by similar vertices thus quantitatively similar expected LDMX signals and relic targets: the first with the kinetic mixing model, the second with the magnetic and electric dipole models, and the third with the charge radius and anapole moment models.
@@ Appendix A Matrix Elements
$\langle\sigma v\rangle_{\chi\overline{\chi}\to f\overline{f}}=\frac{1}{8m_{\chi}^{4}TK_{2}\left(m_{\chi}/T\right)^{2}}\int\sigma(s)_{\chi\overline{\chi}\to f\overline{f}}[s-4m_{\chi}^{2}]\sqrt{s}K_{1}\left(\sqrt{s}/T\right)ds$ (9)
@@
$\sigma_{\chi\overline{\chi}\to f\overline{f}}=\frac{|\boldsymbol{p}_{f}|}{64\pi^{2}s|\boldsymbol{p}_{\chi}|}\int\overline{|\mathcal{M}|^{2}}d\Omega,$ (10)
@@
$\overline{|\mathcal{M}|^{2}}_{\text{KM}}=\frac{2g_{D}^{2}\epsilon^{2}e^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}^{2}})\left[s-(s-4m_{f}^{2})\cos^{2}\theta\right]$ (11)
$\overline{|\mathcal{M}|^{2}}_{\text{M}}=\frac{4sg_{D}^{2}\mu_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})\left[(s-4m_{f}^{2})\cos^{2}\theta+4m_{f}^{2}\right]$ (12)
$\overline{|\mathcal{M}|^{2}}_{\text{E}}=\frac{4sg_{D}^{2}d_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})(s-4m_{f}^{2})\cos^{2}\theta$ (13)
$\overline{|\mathcal{M}|^{2}}_{\text{C}}=\frac{4s^{2}g_{D}^{2}b_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})\left[s-(s-4m_{f}^{2})\cos^{2}\theta\right]$ (14)
$\overline{|\mathcal{M}|^{2}}_{\text{A}}=\frac{4s^{2}g_{D}^{2}a_{f}^{2}}{(s-m_{A^{\prime}}^{2})^{2}+m_{A^{\prime}}^{2}\Gamma_{A^{\prime}}^{2}}(s-4m_{\text{DM}}^{2})(s-4m_{f}^{2})\sin^{2}\theta$ (15)
@@
$\overline{|\mathcal{M}_{A^{\prime}\to\chi\overline{\chi}}|^{2}}=\frac{g_{D}^{2}}{3}(m_{A^{\prime}}^{2}-4m_{\text{DM}}^{2})$ (16)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{KM}}=\frac{e^{2}\epsilon^{2}}{3}(4m_{A^{\prime}}^{2}+8m_{f}^{2})$ (17)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{M}}=\frac{\mu_{f}^{2}}{3}(2m_{A^{\prime}}^{4}+16m_{f}^{2}m_{A^{\prime}}^{2})$ (18)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{E}}=\frac{d_{f}^{2}}{3}(m_{A^{\prime}}^{4}-4m_{f}^{2}m_{A^{\prime}}^{2})$ (19)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{C}}=\frac{b_{f}^{2}}{3}(4m_{A^{\prime}}^{6}+8m_{f}^{2}m_{A^{\prime}}^{4})$ (20)
$\overline{|\mathcal{M}_{A^{\prime}\to f\overline{f}}|^{2}}_{\text{A}}=\frac{a_{f}^{2}}{3}(4m_{A^{\prime}}^{6}-16m_{f}^{2}m_{A^{\prime}}^{4})$ (21)